A rectangular area of 40 square miles has a length equal to three times the width less 2 miles. Find the dimensions of this area.
Answer by checkley77(12844)

L=3W-2
LW=40
(3W-2)W=40
3W^2-2W=40
3W^2-2W-40=0
(3W+10)(W-4)=0
W-4=0
W=4 ANS.
L=3*4-2
L=12-2
L=10 ANS.
PROOF:
4*10=40
40=40